# Mitigating Hallucinations in Large Vision-Language Models with Instruction Contrastive Decoding

### v1

@@ 1 Introduction
Recent research in large vision-language models (LVLMs) Liu et al. (2023c, b); Li et al. (2023a) has seen remarkable progress, benefiting from the integration of advanced large language models (LLMs) Achiam et al. (2023); Touvron et al. (2023a, b) known for their robust language generation and zero-shot transfer capabilities. In order to leverage off-the-shell LLMs, it is crucial to facilitate cross-modal alignment. LLaVa Liu et al. (2023c) employs a linear projection approach, while BLIP-2 Li et al. (2023a) and InstructBLIP Liu et al. (2023b) narrow the modality gap using a Q-Former. Although LVLMs have shown promising outcomes, the issue of hallucination remains. This phenomenon occurs when the generated textual content, despite being fluent and coherent, does not accurately reflect the factual visual content.
@@
The intertwined nature of modalities presents significant challenges in identifying the root causes of hallucinations in LVLMs. Research efforts have begun to uncover the primary contributors to LVLM hallucinations, including statistical biases You et al. (2023) encountered during the training process and excessive dependence on language priors Zhu et al. (2020); Zhibo et al. (2023). Additionally, multimodal misalignment has been identified as a key factor in the occurrence of hallucinations Jiang et al. (2023); Liu et al. (2023a). To address dataset bias, annotation enrichment techniques Gunjal et al. (2024); You et al. (2023); Zhai et al. (2023) have been introduced. Furthermore, to counteract the influence of language priors, post-processing strategies Yin et al. (2023); Zhou et al. (2023) have been developed, along with comprehensive initiatives aimed at improving multimodal alignment through optimizing alignment with humans Sun et al. (2023); Jiang et al. (2023). While these interventions have proven to be effective in reducing hallucinations, they demand substantial human involvement and incur significant computational costs for additional training or the integration of supplementary modules.
In this work, we reveal that appending instructions with role prefixes to form disturbance instructions can significantly exacerbate hallucinations. We hypothesize that identifying and subsequently detaching hallucination concepts from the original distribution could effectively reduce such hallucinations. Motivated by this insight, we introduce the Instruction Contrastive Decoding (ICD) method. This approach is novel in that it is training-free and agnostic to the underlying LVLMs. ICD differentiates between two distributions: one from the original instruction and another from the disturbance instruction within the multimodal alignment module. Utilizing their difference, we aim at suppressing hallucinations. Through comprehensive experiments on discrimination hallucination benchmarks such as POPE Li et al. (2023c) and MME hallucination sets Fu et al. (2023), as well as the generation hallucination benchmark LLaVa-Bench Liu et al. (2023c), our method incorporating state-of-the-art LVLMs like miniGPT4 and InstructBLIP, demonstrates significant efficacy in mitigating hallucinations at both object and attribute levels. Furthermore, our approach consistently enhances performance across general perception and recognition tasks. Our main contributions are as follows:
@@ 2 Related Work
Large Vision-Language Models: The field of vision-language pre-training (VLP) Radford et al. (2021); Li et al. (2022); Bao et al. (2022); Wang et al. (2023a) and fine-tuning Wang et al. (2023b); Wiehe et al. (2022); Alayrac et al. (2022) have seen rapid advancements, propelled by the evolution of large language models (LLMs). As a result, large vision-language models (LVLMs) have emerged, leveraging the strengths of frozen LLMs while emphasizing the facilitating of multimodal alignment modules. Notably, models such as LLaVa and Qwen-VL Bai et al. (2023) adopt simple linear projections to achieve alignment, contrasting with BLIP-2 and miniGPT4 Zhu et al. (2023), which introduce a Q-Former. In further work, InstructBLIP integrates task-aware instructions, enriching the understanding of task-aware visual semantics. Our research builds upon these advancements in LVLMs, focusing on the impact of instruction disturbances. We explore how such disturbances increase the uncertainty in multimodal alignment, significantly contributing to the exacerbation of hallucinations.
Hallucination in VLMs: Hallucination manifests as detailed, fluent, and coherent responses that inaccurately reflect the visual context, including erroneous objects, attributes, and relations Liu et al. (2024). Various strategies have been proposed to curb hallucinations. Annotation enrichment techniques like M-HalDetect Gunjal et al. (2024) and GRIT You et al. (2023), as well as approaches such as HACL Jiang et al. (2023) and LLaVA-RLHL Liu et al. (2023a), seek to improve alignment with human instructions through additional annotations. Similarly, Woodpecker Yin et al. (2023) introduces a post-processing aimed at mitigating biases from language priors. While these methods have shown promise in reducing hallucinations, they often require extensive data annotation, fine-tuning, and supplementary modules, complicating their implementation. In contrast, our method directly addresses hallucinations during inference. Additionally, Leng et al. (2023) introduced a visual contrastive decoding (VCD) approach that contrasts with the distributions of distorted visual inputs, a concept that bears resemblance to our method. However, our ICD method suppresses hallucinations through disturbance instructions affecting multimodal alignment.
@@ 3.1 Inference in LVLMs
Figure 1: An illustration on inference framework and contrastive decoding process of ICD method. At the core (middle orange box), the framework integrates a frozen image encoder, LLM, and query vectors (gray box) within the Q-Former, focusing solely on adjusting the standard and disturbance instructions. The latter, exemplified by adding role prefixes like ’You are a confused object detector,’ aims to increase multimodal alignment uncertainty. This results in two distinct distributions: one from the standard instruction and another influenced by the disturbance. The contrastive decoding method (right orange box) highlights how disturbance instructions amplify hallucinated concepts (’person and fork’), which are then corrected by subtracting probabilities derived from the standard instruction, ensuring accurate recognition of the correct concept ’dog’.
@@
where $y_{<t}$ represents the sequence of generated tokens up to the time steps $(t-1)$. In the decoding phase of LVLMs, hallucinations often emerge when probable tokens lack grounding in the visual context. Jiang et al. (2023); Liu et al. (2023a) indicates that multimodal misalignment is a critical factor contributing to the generation of hallucinations. Thus, we conduct an in-depth analysis of the fusion module, specifically focusing on multimodal alignment. Our work first demonstrates that instructions within the multimodal alignment module can exacerbate hallucinations. To address this, we introduce instruction disturbance and propose an instruction contrastive decoding method, employing a highlight and then detach strategy.
@@ 4.1.2 LVLM Baselines
We employ two state-of-the-art LVLMs as backbone frameworks. Specifically, we implement our ICD on InstructBLIP and miniGPT4, which utilize the Vicuna 7B as their underlying LLM and the sophisticated Q-Former architecture for fusion modules, respectively. Additionally, we explore the use of LLaVa-1.5 Liu et al. (2023b), which incorporates linear projection for its fusion module alongside InstructBLIP, to identify optimal practices in applying the ICD method (see in appendix D). Finally, we compare our method against the visual contrastive decoding approach Leng et al. (2023), designed to mitigate hallucinations arising from visual uncertainties. We posit that our method, being LVLM-agnostic, can be conveniently integrated into various off-the-shelf LVLMs.
@@ 4.3 Discussions on ICD and VCD
This exploration suggests a promising avenue for future research aimed at optimally amalgamating the advantages of both methods. Detailed results and comprehensive analysis of the combined method performance across full MME are provided in the appendix C for further reference.
@@ Appendix B Qualitative Evaluation on LLava-Bench
Appendix C Further Analysis on VCD-Enhanced ICD
@@
Appendix D Optimal Position to Apply Contrastive Decoding

### v2

@@ 1 Introduction
Recent research in large vision-language models (LVLMs) Liu et al. (2023c, b); Li et al. (2023a) has seen remarkable progress, benefiting from the integration of advanced large language models (LLMs) Achiam et al. (2023); Touvron et al. (2023a, b) known for their robust language generation and zero-shot transfer capabilities Zhong et al. (2023); Peng et al. (2023). In order to leverage off-the-shell LLMs, it is crucial to facilitate cross-modal alignment. LLaVa Liu et al. (2023c) employs a linear projection approach, while BLIP-2 Li et al. (2023a) and InstructBLIP Liu et al. (2023b) narrow the modality gap using a Q-Former. Although LVLMs have shown promising outcomes, the issue of hallucination remains. This phenomenon occurs when the generated textual content, despite being fluent and coherent, does not accurately reflect the factual visual content.
@@
The intertwined nature of modalities presents significant challenges in identifying the root causes of hallucinations in LVLMs. Research efforts have begun to uncover the primary contributors to LVLM hallucinations, including statistical biases You et al. (2023) encountered during the training process and excessive dependence on language priors Zhu et al. (2020); Zhibo et al. (2023). Additionally, multimodal misalignment has been identified as a key factor in the occurrence of hallucinations Jiang et al. (2023); Liu et al. (2023a); Wang et al. (2023a). To address dataset bias, annotation enrichment techniques Gunjal et al. (2024); You et al. (2023); Zhai et al. (2023) have been introduced. Furthermore, to counteract the influence of language priors, post-processing strategies Yin et al. (2023); Zhou et al. (2023) have been developed, along with comprehensive initiatives aimed at improving multimodal alignment through optimizing alignment with humans Sun et al. (2023); Jiang et al. (2023). While these interventions have proven to be effective in reducing hallucinations, they demand substantial human involvement and incur significant computational costs for additional training or the integration of supplementary modules.
In this work, we reveal that appending instructions with role prefixes to form disturbance instructions can significantly exacerbate hallucinations. We hypothesize that identifying and subsequently detaching hallucination concepts from the original distribution could effectively reduce such hallucinations. Motivated by this insight, we introduce the Instruction Contrastive Decoding ($\mathcal{\bf ICD}$) method. This approach is novel in that it is training-free and agnostic to the underlying LVLMs. ICD differentiates between two distributions: one from the original instruction and another from the disturbance instruction within the multimodal alignment module. Utilizing their difference, we aim at suppressing hallucinations. Through comprehensive experiments on discrimination hallucination benchmarks such as POPE Li et al. (2023c) and MME hallucination sets Fu et al. (2023), as well as the generation hallucination benchmark LLaVa-Bench Liu et al. (2023c), our method incorporating state-of-the-art LVLMs like miniGPT4 and InstructBLIP, demonstrates significant efficacy in mitigating hallucinations at both object and attribute levels. Furthermore, our approach consistently enhances performance across 14 general perception and recognition tasks within the full MME benchmark.
Our main contributions are as follows:
@@ 2 Related Work
Large Vision-Language Models.
The field of vision-language pre-training (VLP) Radford et al. (2021); Li et al. (2022); Bao et al. (2022); Wang et al. (2023b) and fine-tuning Wang et al. (2023c); Wiehe et al. (2022); Alayrac et al. (2022) have seen rapid advancements, propelled by the evolution of large language models (LLMs). As a result, large vision-language models (LVLMs) have emerged, leveraging the strengths of frozen LLMs while emphasizing the facilitating of multimodal alignment modules. Notably, models such as LLaVa and Qwen-VL Bai et al. (2023) adopt simple linear projections to achieve alignment, contrasting with BLIP-2 and miniGPT4 Zhu et al. (2023), which introduce a Q-Former. In further work, InstructBLIP integrates task-aware instructions, enriching the understanding of task-aware visual semantics. Our research builds upon these advancements in LVLMs, focusing on the impact of instruction disturbances. We explore how such disturbances increase the uncertainty in multimodal alignment, significantly contributing to the exacerbation of hallucinations.
Hallucination in VLMs.
Generation hallucinations have been extensively studied in the field of language modeling Tonmoy et al. (2024); Wen et al. (2024). Hallucination in VLMs manifests as detailed, fluent, and coherent responses that inaccurately reflect the visual context, including erroneous objects, attributes, and relations Liu et al. (2024); Jing et al. (2023). Various strategies have been proposed to curb hallucinations. Annotation enrichment techniques like M-HalDetect Gunjal et al. (2024) and GRIT You et al. (2023), as well as approaches such as HACL Jiang et al. (2023) and LLaVA-RLHL Liu et al. (2023a), seek to improve alignment with human instructions through additional annotations. Similarly, Woodpecker Yin et al. (2023) introduces post-processing aimed at mitigating biases from language priors. While these methods have shown promise in reducing hallucinations, they often require extensive data annotation, fine-tuning, and supplementary modules, complicating their implementation. In contrast, our method directly addresses hallucinations during inference. Additionally, Leng et al. (2023) introduced a visual contrastive decoding (VCD) approach that contrasts with the distributions of distorted visual inputs, a concept that bears resemblance to our method. However, our ICD method suppresses hallucinations through disturbance instructions affecting multimodal alignment.
@@ 3.1 Inference in LVLMs
Figure 1: An illustration on inference framework and contrastive decoding process of ICD method. At the core (middle orange box), the framework integrates a frozen image encoder, LLM, and query vectors (gray box) within the Q-Former, focusing solely on adjusting the standard and disturbance instructions. The latter, exemplified by adding role prefixes like ‘You are a confused object detector,’ aims to increase multimodal alignment uncertainty. This results in two distinct distributions: one from the standard instruction and another influenced by the disturbance. The contrastive decoding method (right orange box) highlights how disturbance instructions amplify hallucinated concepts (‘person and fork’), which are then corrected by subtracting probabilities derived from the standard instruction, ensuring accurate recognition of the correct concept ‘dog’.
@@
where $y_{<t}$ represents the sequence of generated tokens up to the time steps $(t-1)$. In the decoding phase of LVLMs, hallucinations often emerge when probable tokens lack grounding in the visual context. Jiang et al. (2023); Liu et al. (2023a) indicate that multimodal misalignment is a critical factor contributing to the generation of hallucinations. Thus, we conduct an in-depth analysis of the fusion module, specifically focusing on multimodal alignment. Our work first demonstrates that instructions within the multimodal alignment module can exacerbate hallucinations. To address this, we introduce instruction disturbance and propose an instruction contrastive decoding method, employing a highlight and then detach strategy.
@@ 4.1.2 LVLM Baselines
We employ two state-of-the-art LVLMs as backbone frameworks. Specifically, we implement our ICD on InstructBLIP and miniGPT4, which utilize the Vicuna 7B as their underlying LLM and the sophisticated Q-Former architecture for fusion modules, respectively. Additionally, we explore the use of LLaVa-1.5 Liu et al. (2023b), which incorporates linear projection for its fusion module alongside InstructBLIP, to identify optimal practices in applying the ICD method (see in appendix E). Finally, we compare our method against the visual contrastive decoding approach Leng et al. (2023), designed to mitigate hallucinations arising from visual uncertainties. We posit that our method, being LVLM-agnostic, can be conveniently integrated into various off-the-shelf LVLMs.
@@ 4.3 Discussions on ICD and VCD
This exploration suggests a promising avenue for future research aimed at optimally amalgamating the advantages of both methods. Detailed results and comprehensive analysis of the combined method performance across full MME are provided in the appendix D for further reference.
@@ 5 Conclusion
Acknowledgments
We thank the anonymous reviewers and the area chair for their insightful comments and suggestions. This research was funded by the German Research Foundation DFG Transregio SFB 169: Crossmodal Learning: Adaptivity, Prediction, and Interaction.
@@ Appendix A Implementation Details
We conducted experiments with various instructional disturbances, incorporating both positive and negative role prefixes. For illustrative purposes, we present two positive role prefix instructions and two negative role prefix instructions, providing a detailed guide for others to effectively implement our method in practical applications.
  1. $\square$
    P1: You are an object detector to recognize every different object.
  2. $\square$
    P2: You are an object detector to recognize every different object by focusing on the shapes, colors, and relationships of objects.
  3. $\triangle$
    N1: I want you to avoid any specific identification or categorization of the objects depicted.
  4. $\triangle$
    N2: You are a confused object detector to provide a fuzzy overview or impression of the image.
@@ Appendix B Qualitative Evaluation on LLaVa-Bench
Appendix C General QA Benchmarks Performance
Our evaluation in Section 4, utilizing the POPE, MME, and LLava-Bench benchmarks, focused on assessing hallucinations in LVLMs, particularly regarding the presence of objects and attributes within images. In this section, we broaden our evaluation scope to include more general QA and caption datasets, such as MSCOCO Lin et al. (2014), OK-VQA Marino et al. (2019), and TextVQA Singh et al. (2019), using metrics like CHAIR Rohrbach et al. (2018), CIDEr Vedantam et al. (2015), and BLEU Papineni et al. (2002) for a comprehensive analysis.
Specifically, MSCOCO comprises 118,000 images in the training set and 5,000 images in the validation set. Following the evaluation settings of Rohrbach et al. (2018), we report instance-level (CHAIR_I) and sentence-level (CHAIR_S) hallucination results using the MSCOCO validation set in Table  3. Additionally, in line with Yue et al. (2024), we randomly sample 500 instances from the MSCOCO training and validation sets, with our results presented in Table  4.
Method CHAIR_I $\downarrow$ CHAIR_S $\downarrow$
InstructBLIP 10.7 20.0
VCD 9.3 18.2
ICD 8.0 15.2
Table 3: Evaluation on MSCOCO validation set using metrics CHAIR_I and CHAIR_S, instance-level and sentence-level hallucinations, followed by Rohrbach et al. (2018).
Method CHAIR_I $\downarrow$ CHAIR_S $\downarrow$
InstructBLIP 12.2 21.4
VCD 9.0 16.6
ICD 7.7 14.4
Table 4: Evaluation on MSCOCO training and validation sets (500 samples), using metrics CHAIR_I and CHAIR_S, instance-level and sentence-level hallucinations, followed by Yue et al. (2024).
We further evaluate our ICD method on the OK-VQA and TextVQA datasets using CIDEr and BLEU metrics, with the improved results detailed in Tables  5 and  6.
Method CIDEr BLEU1 BLEU2
InstructBLIP 0.28 0.33 0.17
VCD 0.35 0.42 0.22
ICD 0.40 0.45 0.25
Table 5: Evaluation on OK-VQA test set using metrics CIDEr and BLEU 1 and 2.
Method CIDEr BLEU1 BLEU2 BLEU3 BLEU4
InstructBLIP 0.56 0.29 0.22 0.19 0.19
VCD 0.71 0.36 0.32 0.30 0.31
ICD 0.69 0.35 0.30 0.29 0.30
Table 6: Evaluation on TextVQA test set using metrics CIDEer and BLEU 1, 2, 3, and 4.
These additional evaluations underscore the versatility and efficacy of the ICD method across diverse QA datasets and evaluation metrics. Notably, the improvements in metrics such as CHAIR, CIDEr, and BLEU across the MSCOCO, OK-VQA, and TextVQA benchmarks reaffirm the significant impact of our method.
Appendix D Further Analysis on VCD-Enhanced ICD
@@
Appendix E Optimal Position to Apply Contrastive Decoding